# Vacuum Rabi Splitting and Quantum Fisher Information of a Non-Hermitian Qubit in a Single-Mode Cavity

### v1

@@ Abstract
A natural extension of the non-Hermitian qubit is to place it in a single-mode cavity. This setup corresponds to the quantum Rabi model (QRM) with a purely imaginary bias on the qubit, exhibiting the parity-time ($\mathcal{P}\mathcal{T}$) symmetry. In this work, we first solve the $\mathcal{P}\mathcal{T}$ symmetric QRM using the Bogoliubov operator approach. We derive the transcendental function responsible for the exact solution, which can also be used to precisely identify exceptional points. The adiabatic approximation previously used can be easily formulated within this approach by considering transitions between the same manifolds in the space of Bogoliubov operators. By further considering transitions between the nearest-neighboring manifolds, we can analytically obtain more accurate eigen-solutions. Moreover, these simple corrections can capture the main features of the dynamics, where the adiabatic approximation fails. Furthermore, the rich characteristics of the vacuum Rabi splitting in the emission spectrum is predicted. The width of the peaks increases with the coupling strength and the imaginary biases, reflecting the nature of open quantum systems. Additionally, we identify a quantum criticality-enhanced effect by calculating the quantum Fisher information. Near the exceptional points, the quantum Fisher information in the $\mathcal{P}\mathcal{T}$ symmetric QRM is significantly higher than that of the non-Hermitian qubit component. This may open a new avenue for enhancing quantum sensitivity in non-Hermitian systems by incorporating coupling with an additional degree of freedom, enabling more precise parameter estimation.
@@ I Introduction
In recent years, concepts from non-Hermitian physics and parity-time ($\mathcal{P}$ $\mathcal{T}$)-symmetry have attracted considerable interest. Bender et al. demonstrated that a non-Hermitian Hamiltonian invariant under combined space reflection and time reversal can exhibit a real spectrum [1]. This counterintuitive result challenges the traditional view that real eigenvalues are associated exclusively with Hermitian observables, marking a significant milestone in non-Hermitian physics [2, 3, 4, 5, 6].
@@
Unlike the strict mathematical requirement in traditional quantum physics that a Hamiltonian operator should be Hermitian, the $\mathcal{PT}$ invariance condition is a weaker, more physically intuitive constraint, providing a rich field for both theoretical and experimental exploration. As system parameters varying, the $\mathcal{PT}$-symmetric system can transition between the $\mathcal{PT}$-unbroken and $\mathcal{P}\mathcal{T}$-broken phases. In the $\mathcal{PT}$-unbroken phase, the system has real eigenvalues, while in the $\mathcal{PT}$-broken phase, it has a pair of conjugate eigenvalues. Phase transitions are characterized by numerous accumulation points of exceptional points (EPs). At these EPs, both eigenvectors and eigenvalues coalesce. Additionally, dimensionality reduction occurs in the eigenspace of the EPs [7, 8].
@@
Non-Hermitian $\mathcal{PT}$-symmetry has recently received considerable attention in the study of $\mathcal{PT}$-symmetric systems involving the interaction of qubits with classical light fields. [9, 10, 11, 9, 12]. This research belongs to the context of the non-Hermitian semi-classical Rabi model. $\mathcal{PT}$-symmetry in pure quantum systems, particularly in the context of light-matter interaction with an imaginary qubit-cavity coupling, has also been studied. Examples include the non-Hermitian Jaynes-Cummings model [13], the non-Hermitian double Jaynes-Cummings model [14], the dissipative quantum Rabi model (QRM) [15], and the non-Hermitian QRM [16].
@@
The $\mathcal{PT}$-symmetric quantum Rabi model (PTQRM) with a purely imaginary bias [17] has rarely been explored. This model is, however, effectively a natural extension of the well-studied non-Hermitian two-level system (NH-TLS) coupled to a quantized light field. The PTQRM exhibits intriguing non-Hermitian properties, including an infinite number of EPs that vanish and reappear depending on the strength of the light-matter coupling. Furthermore, the PTQRM can be realized in realistic physical systems [17]. Potential physical implementations of the non-Hermitian Hamiltonian for the PTQRM include circuit quantum electrodynamics (cQED) [18], trapped-ion experiments, and the replacement of traditional qubits in QRM setups [18, 19, 20, 21, 22].
@@
In this work, we extend the Bogoliubov operator approach (BOA) [23, 24, 25] to solve the PTQRM [17], and obtain the exact eigen-solutions and the EPs. Within this framework, we analytically propose a correction to the previous adiabatic approximation (AA), especially the eigenstates, leading to a deeper understanding of dynamics. We primarily investigate two prominent quantum effects of this model using our developed approaches: vacuum Rabi splitting (VRS), a key phenomenon in atomic physics  [26], and Quantum Fisher information (QFI), which plays a crucial role in quantum metrology. Recent proposals suggest that quantum sensors based on non-Hermitian systems may exhibit enhanced sensitivity  [27, 28, 29, 30, 31, 32, 33, 34, 35]. We calculate the QFI for both the ground and excited states of the system, with a particular focus on the QFI near the EPs [29, 15].
@@
This paper is organized as follows: In Sec. II, we derive the transcendental function whose zeros provide the complete spectrum, using the Bogoliubov operator approach (BOA). The EPs are identified through their real number form in the $\mathcal{PT}$-unbroken phase. The previously often used AA is further corrected analytically. In Sec. III, we examine the atomic dynamics and the VRS in the emission spectrum using several approaches. The novel characteristics of the VRS are explored and discussed in terms of analytical eigenstates. In Sec. IV, we calculate the QFI in the PTQRM and its Hermitian counterpart, as well as for the NH-TLS, and perform the comparison of the QFI values near the EPs. Finally, a brief summary is provided in Sec. V. The appendix presents the determinant of the CAA and the selection rules for the eigenstates.
@@ II Analytical solution
where the first two terms represent a non-Hermitian qubit with energy splitting $\Delta$ and an imaginary bias $i\epsilon$, $\sigma_{x,z}$ are the Pauli matrices, $a^{{\dagger}}$ and $a$ denote the creation and annihilation operators with cavity frequency $\omega$, and $g$ represents the qubit-cavity coupling strength. Its Hermitian counterpart is just the asymmetric QRM [36, 23, 37, 38, 39, 40, 41, 42] For $\epsilon=0$, the Hamiltonian (1) exhibits $\mathbb{Z}_{2}$ symmetry [43]. The conserved parity operator is defined as $\hat{P}=-\sigma_{x}\exp(i\pi a^{{\dagger}}a)$, with eigenvalues ±${\pm 1}$. This symmetry allows the Hamiltonian (1) to be diagonalized into two blocks, each corresponding to a different parity subspace. Due to this $\mathbb{Z}_{2}$ symmetry, eigenvalues from different symmetry parity are allowed to cross.
When a pure imaginary bias $\epsilon\neq 0$ is introduced, the parity operator $\hat{P_{1}}$ is no longer conserved. Combined with the standard time-reversal operator $T$, which takes the complex conjugate, and noting that $\mathcal{T}^{\dagger}\hat{x}\mathcal{T}=\hat{x}$ and $\mathcal{T}^{\dagger}\hat{p}\mathcal{T}=-\hat{p}$, where $\hat{x}$ is the displacement operator and $\hat{p}$ is the momentum operator, it can be verified that $\left[H,\mathcal{PT}\right]=0$. Therefore, the non-Hermitian Hamiltonian (1) is indeed $\mathcal{PT}$-symmetric, and is thus termed the PTQRM.
The PTQRM can be implemented in a practical system as proposed in Ref. [17]. After performing a rotation around the $y$-axis in the atomic sector, the Hamiltonian (1) transforms into their Eq. (1). Consider a passive $\mathcal{PT}$-symmetric qubit with the Hamiltonian $H_{q}^{p}=-\frac{\Delta}{2}\sigma_{x}+i\epsilon\sigma_{+}\sigma_{-}$, which differs from the standard $\mathcal{PT}$-symmetric qubit by a constant shift of $i\epsilon/2$, but retains all the essential features of standard $\mathcal{PT}$ symmetry, including EPs. Incorporating a cavity into this passive qubit leads to the QRM under passive $\mathcal{PT}$-symmetry, known as the passive PTQRM. Using the Lindblad master-equation approach, Lu et al. demonstrate that the effective Hamiltonian governing the dynamics is synonymous with the passive PTQRM in the rotating frame, and is equivalent to the standard PTQRM Hamiltonian  [17]. These authors have also proposed a potential circuit QED realization in which the transmon qubit circuit is embedded in a cavity.
@@
II.1 Exact Solutions by the BOA
@@
the Hamiltonian (1) can be transformed into the following matrix form with units $\hbar=\omega=1$.
$H=\left(\begin{array}[]{cc}A_{+}^{\dagger}A_{+}+i\frac{\epsilon}{2}-g^{2}&-\Delta/2\\ -\Delta/2&A_{-}^{\dagger}A_{-}-i\frac{\epsilon}{2}-g^{2}\end{array}\right),$ (3)
@@
Figure 1: ln$\lvert G\rvert^{2}$, where G is calculated from the G-function (10), in the complex $E$ plane. Open circles mark the zeros of both the real and imaginary parts of the G-function, which coincide exactly with the eigenvalues from the ED. The parameters are $\Delta=0.5$, $\epsilon=0.2$, and $g=0.25$.
Figure 2: (colored online) (a) G-curves from Eq. (11) (blue line) at $g=0.1$ in the real eigenvalue regime, with circles representing ED results. (b) The real part of eigenvalues as a function of $g$ from the ED, with the dashed lines marking the real part of eigenvalues in the $\mathcal{P}\mathcal{T}$-broken region. Crossing values between the vertical dotted lines at $g=0.1,0.5$ and $0.6828$ and real eigenvalues (solid line) in the $\mathcal{P}\mathcal{T}$-unbroken region are confirmed in the G-functions in (a) and the lower panel. G-curves from Eq. (11) at $g=0.5,0.6828$ and $0.7$ are shown in (c), (d), and (e) with blue lines, respectively. The parameters are $\Delta=0.5$ and $\epsilon=0.2$.
If both wavefunctions (4) and (5) are true eigenfunctions for a non-degenerate eigenstate with eigenvalue $E$, they should differ only by a constant ratio, i.e. $\left|A_{+}\right\rangle=z\left|A_{-}\right\rangle$. Projecting both sides onto the original vacuum state $\left|0\right\rangle$, using $\sqrt{n!}{\langle}0|n{\rangle}_{A_{+}}=(-1)^{n}\sqrt{n!}{\langle}0|n{\rangle}_{A_{-}}=e^{-g^{2}/2}g^{n}$ and eliminating $z$ gives
@@
Using Eqs. (7) and (8), one can derive the complex G-function of the PTQRM,
@@
×$\times\left[\sum_{n=0}^{\infty}\frac{c_{n}}{n-g^{2}-\frac{i\epsilon}{2}-E}g^{n}\right]$
@@
It includes both real and imaginary components. The zeros of this G-function provide both the real and imaginary components of the eigenvalues. As illustrated in Fig. 1, the eigenvalues obtained from the G-function (10) show excellent agreement with those from exact diagonalization (ED).
@@
By following the same procedure, one can derive the real G-function in the form
@@
Although the coefficients $e_{n}$ and $f_{n}$ are complex, both terms in Eq. (11) represent squares of the modulus, making the G-function real. Its zeros can determine all real eigenvalues of this model within the $\mathcal{PT}$-unbroken region.
As shown in Fig. 2, all real eigenvalues (solid line) can be obtained by zeros of this real G-functions. EPs occur when both the G-function in Eq. (11) and its first-order derivative with respect to E are simultaneously zero, as illustrated in Fig. 2 (d). The zero of the first-order derivative identifies the EP where two real eigenvalues merge into one. Thus, using the real G-function Eq. (11), we can accurately pinpoint the position of the EPs and show that both eigenstates and eigenvalues coalesce at these points. It should be noted that, while we benefit from using the real G-function (11) to determine the EPs, it cannot be applied to $\mathcal{PT}$-broken region, marked by the dotted line in Fig. 2 (b) and the none solution in panel (e).
Finally, from Eqs. (7) and (8), it is clear that the denominator cannot vanish for any E due to the presence of the imaginary bias $\epsilon$. As a result, the pole structure of the G-functions does not exist, in sharp contrast to its Hermitian counterpart  [36, 23]. In other words, all eigenvalues can be given by the zeros of the general G-function (10), and all real eigenvalues in the $\mathcal{PT}$unbroken phase can be obtained by zeros of the real G-function Eq.(11), even at the EPs, in the PTQRM. Note that the doubly degenerate solutions in its Hermitian counterpart cannot be given by the corresponding G-function  [38, 25].

### v2

@@ Abstract
A natural extension of the non-Hermitian qubit is to place it in a single-mode cavity. This setup corresponds to the quantum Rabi model (QRM) with a purely imaginary bias on the qubit, exhibiting parity-time ($\mathcal{P}\mathcal{T}$) symmetry. In this work, we first solve the $\mathcal{P}\mathcal{T}$-symmetric QRM using the Bogoliubov operator approach. We derive the transcendental function responsible for the exact solution, which can also be used to precisely identify exceptional points. The adiabatic approximation previously used can be easily formulated within this approach by considering transitions between the same manifolds in the space of Bogoliubov operators. By further considering transitions between the nearest-neighboring manifolds, we can analytically obtain more accurate eigensolutions. Moreover, these simple corrections can capture the main features of the dynamics, where the adiabatic approximation fails. Furthermore, the rich characteristics of the vacuum Rabi splitting in the emission spectrum are predicted. The width of the peaks increases with the coupling strength and the imaginary biases, reflecting the nature of open quantum systems. Additionally, we identify a quantum-criticality-enhanced effect by calculating the quantum Fisher information. Near the exceptional points, the quantum Fisher information in the $\mathcal{P}\mathcal{T}$-symmetric QRM is significantly higher than that of the non-Hermitian qubit component. This may open a new avenue for enhancing quantum sensitivity in non-Hermitian systems by incorporating coupling with an additional degree of freedom, enabling more precise parameter estimation.
@@ I Introduction
In recent years, concepts from non-Hermitian physics and parity-time ($\mathcal{P}$$\mathcal{T}$) symmetry have attracted considerable interest. Bender et al. demonstrated that a non-Hermitian Hamiltonian invariant under combined space reflection and time reversal can exhibit a real spectrum [1]. This counterintuitive result challenges the traditional view that real eigenvalues are associated exclusively with Hermitian observables, marking a significant milestone in non-Hermitian physics [2, 3, 4, 5, 6].
@@
Unlike the strict mathematical requirement in traditional quantum physics that a Hamiltonian operator should be Hermitian, the $\mathcal{PT}$ invariance condition is a weaker, more physically intuitive constraint, providing a rich field for both theoretical and experimental exploration. As system parameters vary, the $\mathcal{PT}$-symmetric system can transition between the $\mathcal{PT}$-unbroken and $\mathcal{P}\mathcal{T}$-broken phases. In the $\mathcal{PT}$-unbroken phase, the system has real eigenvalues, while in the $\mathcal{PT}$-broken phase, it has a pair of conjugate eigenvalues. Phase transitions are characterized by numerous accumulation points of exceptional points (EPs). At these EPs, both eigenvectors and eigenvalues coalesce. Additionally, dimensionality reduction occurs in the eigenspace of the EPs [7, 8].
@@
Non-Hermitian $\mathcal{PT}$ symmetry has recently received considerable attention in the study of $\mathcal{PT}$-symmetric systems involving the interaction of qubits with classical light fields [9, 10, 11, 9, 12]. This research belongs to the context of the non-Hermitian semiclassical Rabi model. $\mathcal{PT}$symmetry in pure quantum systems, particularly in the context of light-matter interaction with an imaginary qubit-cavity coupling, has also been studied. Examples include the non-Hermitian Jaynes-Cummings model [13], the non-Hermitian double Jaynes-Cummings model [14], the dissipative quantum Rabi model (QRM) [15], and the non-Hermitian QRM [16].
@@
The $\mathcal{PT}$-symmetric quantum Rabi model (PTQRM) with a purely imaginary bias [17] has rarely been explored. This model is, however, effectively a natural extension of the well-studied non-Hermitian two-level system (NHTLS) coupled to a quantized light field. The PTQRM exhibits intriguing non-Hermitian properties, including an infinite number of EPs that vanish and reappear depending on the strength of the light-matter coupling. Furthermore, the PTQRM can be realized in realistic physical systems [17]. Potential physical implementations of the non-Hermitian Hamiltonian for the PTQRM include circuit quantum electrodynamics (cQED) [18], trapped-ion experiments, and the replacement of traditional qubits in QRM setups [18, 19, 20, 21, 22].
@@
In this work, we extend the Bogoliubov operator approach (BOA) [23, 24, 25] to solve the PTQRM [17], and obtain the exact eigensolutions and the EPs. Within this framework, we analytically propose a correction to the previous adiabatic approximation (AA), especially the eigenstates, leading to a deeper understanding of dynamics. We primarily investigate two prominent quantum effects of this model using our developed approaches: vacuum Rabi splitting (VRS), a key phenomenon in atomic physics  [26], and Quantum Fisher information (QFI), which plays a crucial role in quantum metrology. Recent proposals suggest that quantum sensors based on non-Hermitian systems may exhibit enhanced sensitivity  [27, 28, 29, 30, 31, 32, 33, 34, 35]. We calculate the QFI for both the ground and excited states of the system, with a particular focus on the QFI near the EPs [29, 15].
@@
This paper is organized as follows: In Sec. II, we derive the transcendental function whose zeros provide the complete spectrum, using the BOA. The EPs are identified through their real number form in the $\mathcal{PT}$-unbroken phase. The previously often used AA is further corrected analytically. In Sec. III, we examine the atomic dynamics and the VRS in the emission spectrum using several approaches. The novel characteristics of the VRS are explored and discussed in terms of analytical eigenstates. In Sec. IV, we calculate the QFI in the PTQRM and its Hermitian counterpart, as well as for the NHTLS, and perform the comparison of the QFI values near the EPs. Finally, a brief summary is provided in Sec. V. The appendix presents the determinant of the symmetric correction to the AA (the CAA) and the selection rules for the eigenstates.
@@ II Analytical solution
where the first two terms represent a non-Hermitian qubit with energy splitting $\Delta$ and an imaginary bias $i\epsilon$, $\sigma_{x,z}$ are the Pauli matrices, $a^{{\dagger}}$ and $a$ denote the creation and annihilation operators with cavity frequency $\omega$, and $g$ represents the qubit-cavity coupling strength. Its Hermitian counterpart is just the asymmetric QRM [36, 23, 37, 38, 39, 40, 41, 42]. For $\epsilon=0$, Hamiltonian (1) exhibits $\mathbb{Z}_{2}$ symmetry [43,44]. The conserved parity operator is defined as $\hat{P}=-\sigma_{x}\exp(i\pi a^{{\dagger}}a)$, with eigenvalues $\pm 1$}. This symmetry allows Hamiltonian (1) to be diagonalized into two blocks, each corresponding to a different parity subspace. Due to this $\mathbb{Z}_{2}$ symmetry, eigenvalues from different symmetry parity are allowed to cross.
When a pure imaginary bias $\epsilon\neq 0$ is introduced, the parity operator $\hat{P}$ is no longer conserved. Combined with the standard time-reversal operator $T$, which takes the complex conjugate, and noting that $\mathcal{T}^{\dagger}\hat{x}\mathcal{T}=\hat{x}$ and $\mathcal{T}^{\dagger}\hat{p}\mathcal{T}=-\hat{p}$, where $\hat{x}$ is the displacement operator and $\hat{p}$ is the momentum operator, it can be verified that $\left[H,\mathcal{PT}\right]=0$. Therefore, the non-Hermitian Hamiltonian ( 1) is indeed $\mathcal{PT}$-symmetric, and is thus termed the PTQRM. The PTQRM can be implemented in a practical system as proposed in Ref. [17]. After performing a rotation around the y axis in the atomic sector, Hamiltonian ( 1) transforms into their Eq. (1). Consider a passive $\mathcal{PT}$-symmetric qubit with the Hamiltonian $H_{q}^{p}=-\frac{\Delta}{2}\sigma_{x}+i\epsilon\sigma_{+}\sigma_{-}$, which differs from the standard $\mathcal{PT}$-symmetric qubit [i.e., first two terms in ( 1)] by a constant shift of $i\epsilon/2$, but retains all the essential features of standard $\mathcal{PT}$ symmetry, including EPs. Incorporating a cavity into this passive qubit leads to the QRM under passive $\mathcal{PT}$ symmetry, known as the passive PTQRM. Using the Lindblad master-equation(LME) approach, Lu et al. demonstrate that the effective Hamiltonian governing the dynamics is synonymous with the passive PTQRM in the rotating frame, and is equivalent to the standard PTQRM Hamiltonian  [17]. These authors have also proposed a potential circuit QED realization in which the transmon qubit circuit is embedded in a cavity.
The effective non-Hermitian Hamiltonian typically emerges within the framework of quantum trajectories and postselection. Reference [45] offers a detailed explanation of when quantum jumps can be neglected and clarifies the relationship between the master equation and the non-Hermitian approach. Discarding quantum jumps is as phenomenological as introducing a complex energy shift. In the absence of quantum jumps, the system evolves according to the effective non-Hermitian Hamiltonian. Repeating experiments and discarding trajectories have been experimentally demonstrated in Ref.  [18], where an effective non-Hermitian Hamiltonian and the EPs for quantum systems were reported. This suggests that experimentalists have already considered the effects of postselection in practical experiments.
@@
II.1 Exact solutions by the Bogoliubov operator approach
@@
Hamiltonian (1) can be transformed into the following matrix form with units $\hbar=\omega=1$:
$H=\left(\begin{array}[]{cc}A_{+}^{\dagger}A_{+}+i\frac{\epsilon}{2}-g^{2}&-\Delta/2\\ -\Delta/2&A_{-}^{\dagger}A_{-}-i\frac{\epsilon}{2}-g^{2}\end{array}\right).$ (3)
@@
[…]
If both wavefunctions (4) and (5) are true eigenfunctions for a nondegenerate eigenstate with eigenvalue $E$, they should differ only by a constant ratio, i.e. $\left|A_{+}\right\rangle=z\left|A_{-}\right\rangle$. Projecting both sides onto the original vacuum state $\left|0\right\rangle$, using $\sqrt{n!}{\langle}0|n{\rangle}_{A_{+}}=(-1)^{n}\sqrt{n!}{\langle}0|n{\rangle}_{A_{-}}=e^{-g^{2}/2}g^{n}$ and eliminating $z$ gives
@@
Thus, the $G$ function can be defined as
$G=\sum_{n=0}^{\infty}e_{n}g^{n}\sum_{n=0}^{\infty}d_{n}g^{n}-\sum_{n=0}^{\infty}f_{n}g^{n}\sum_{n=0}^{\infty}c_{n}g^{n}.$ (10)
By substituting Eqs. (7) and (8), one can arrive at
@@
$\times\left[\sum_{n=0}^{\infty}\frac{c_{n}}{n-g^{2}-\frac{i\epsilon}{2}-E}g^{n}\right]$
@@
It includes both real and imaginary components. The zeros of this $G$ function provide both the real and imaginary components of the eigenvalues. As illustrated in Fig. 1, the eigenvalues obtained from the $G$ function (11) show excellent agreement with those from exact diagonalization (ED).
[…]
@@
By following the same procedure, one can derive the real $G$ function in the form
@@
Although the coefficients $e_{n}$ and $f_{n}$ are complex, both terms in Eq. (12) represent squares of the modulus, making the $G$ function real. Its zeros can determine all real eigenvalues of this model within the $\mathcal{PT}$-unbroken region.
As shown in Fig. 2, all real eigenvalues (solid line) can be obtained by zeros of this real $G$ function. EPs occur when both the $G$ function in Eq. (12) and its first-order derivative with respect to E are simultaneously zero, as illustrated in Fig. 2 (d). The zero of the first-order derivative identifies the EP where two real eigenvalues merge into one. Thus, using the real $G$ function in Eq. (12), we can accurately pinpoint the position of the EPs and show that both eigenstates and eigenvalues coalesce at these points. It should be noted that, while we benefit from using the real $G$ function (12) to determine the EPs, it cannot be applied to $\mathcal{PT}$-broken region, marked by the dotted line in Fig. 2 (b) and the nonsolution in Fig. 2 (e).
Finally, from Eqs. (7) and (8), it is clear that the denominator cannot vanish for any $E$ due to the presence of the imaginary bias $\epsilon$. As a result, the pole structure of the $G$ functions does not exist, in sharp contrast to its Hermitian counterpart [23,26]. In other words, all eigenvalues can be given by the zeros of the general $G$ function (11), and all real eigenvalues in the $\mathcal{PT}$ -unbroken phase can be obtained by zeros of the real $G$ function Eq.(12), even at the EPs, in the PTQRM. Note that the doubly degenerate solutions in its Hermitian counterpart cannot be given by the corresponding $G$ function [25,28].